OpenStudy (anonymous):

plz answer 5th grade math question

DivineSolar (divinesolar):

The image is an X

DivineSolar (divinesolar):

Try using gyazo to screenshot.

OpenStudy (xxstudyingstuffxx):

If you have a windows you can press the windows button + PrtScr. PrtScr is normally above the backspace.

OpenStudy (mathstudent55):

Why does each choice have a "A." in the beginning?

OpenStudy (mathstudent55):

@hackerlife120 Do you still need help with this problem?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

methinks A = Amount A.x = amount of "x" A.y = amount of "y"

OpenStudy (mathstudent55):

Ok. 1 lb of spinach sells for $3.50.

OpenStudy (xxstudyingstuffxx):

Well, if you have learned ratios then you know that there is 1 pound of spinach for every 3.5 dollars. The total amount of dollars would be however many pounds of spinach was bought times the amount of money per pound.

OpenStudy (mathstudent55):

You can use the price of the spinach to figure out how much 2 lb, or 3 lb, or other amounts would cost.

OpenStudy (xxstudyingstuffxx):

Does that make sense?

OpenStudy (anonymous):

no

OpenStudy (mathstudent55):

1 lb of spinach costs $3.50 2 lbs is 2 * 1 lb, so let's multiply both the amount and the price by 2 to see how much 2 lb would cost. 1 lb costs $3.50 2 * 1 lb = 2 lb costs $3.50 * 2 = $7.00 2 lb of spinach costs $7.00 Ok so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im thinking the answer is c

OpenStudy (mathstudent55):

Now let's try 3 lb. Similar to what we did before, 3 lb = 3 * 1 lb. 1 lb costs $3.50 3 * 1 lb = 3 lb costs $3.50 * 3 = $10.50 3 lb costs $10.50

OpenStudy (xxstudyingstuffxx):

y = total amount of money. x = amount of spinach bought. The amount of spinach bought determines how much money is spent. y(the amount of money) = 3.50 * x(the amount of spinach bought.

OpenStudy (anonymous):

cant give both of you medals so give each other one

OpenStudy (anonymous):

soooo b

OpenStudy (xxstudyingstuffxx):

I didn't get one, but I gave mathstudent55 one.

OpenStudy (mathstudent55):

Now instead of trying to find the price of a specific number of pounds, we use the letter x, called a variable in math, to represent any number we want. 1 lb of spinach costs $3.50 How much do x ponds of spinach cost? x = x * 1, so x * 1 lb = x lb of spinach costs $3.50 * x We can simplify the writing and write: x pounds of spinach cost 3.50x

OpenStudy (xxstudyingstuffxx):

And yes, you are correct.

OpenStudy (xxstudyingstuffxx):

Can I get a metal?

OpenStudy (mathstudent55):

Your problem states that x is the number of pounds, and y is the cost. Since x pounds of spinach cost 3.50x, and the cost is y, we get y = 3.50x

OpenStudy (xxstudyingstuffxx):

Take the one you gave him away and give it to me. I gave mine to him.

OpenStudy (xxstudyingstuffxx):

Thx

OpenStudy (anonymous):

@XxStudyingStuffxX

OpenStudy (xxstudyingstuffxx):

@hackerlife120 Same concept as the last question. C, or the amount of money total, is equal to $10.98 times the amount of pounds the customer bought. So, what would the answer be?